OpenStudy (anonymous):

Solve: 4(x + 1) + 8 = 24

OpenStudy (anonymous):

Distribute the 4: 4(x+1)+8=24 4x+4+8=24 Can you solve it now? :o)

OpenStudy (anonymous):

nop

OpenStudy (anonymous):

Alright, well now you're going to combine the like terms of +4 and +8. 4x+12=24 Now, we need to get x alone, so, you're going to get the 12 over to the 24. Can you do it?

OpenStudy (anonymous):

give me another hit to solve it

OpenStudy (anonymous):

Well, to cancel the +12, or to get it's value to 0, you need to do the opposite of what its doing. So, +12 is positive and adding, so you need to the opposite operation of adding and a number to get the +12 to cancel on the left side. And whatever you do to one side you must do to the other. So, whatever you do to the +12 you must do to the 24 on the right side. :o)

OpenStudy (anonymous):

4(x + 1) + 8 = 24 4x+4+8=24 4x+12=24 4x=12 x=3

OpenStudy (anonymous):

Awesome job :o)